write an equation of a vertical line through (4,3).
write an equation of the line parallel to the line 8x - 2y = 7 passing through (2,-1).
find all values of x for which (x,6) is on the graph of f (x) = x2 + 3x -4

Answer by mananth(16946)

write an equation of a vertical line through (4,3).
x=4
write an equation of the line parallel to the line 8x - 2y = 7 passing through (2,-1).
8 x -2 y = 7
Find the slope of this line
make y the subject
-2 y = -8 x + 7
Divide by -2
y = 4 x -3 1/2
Compare this equation with y=mx+b
slope m = 4
The slope of a line parallel to the above line will be the same
The slope of the required line will be 4
m= 4 ,point ( 2 , -1 )
Find b by plugging the values of m & the point in
y=mx+b
-1 = 8 + b
b= 9
m= 4
Plug value of the slope and b in y = mx +b
The required equation isy = 4 x + 9



find all values of x for which (x,6) is on the graph of f (x) = x2 + 3x -4
6=x^2+3x-5
x^2+3x-10=0
(x+5)(x-2)=0
x=-5 & 2